OpenStudy (anonymous):

Solve: 3(x + 4) greater than or equal to x - 2 x greater than or equal to -7 x greater than or equal to 7 x less than or greater to -7 x greater than or equal to -3

OpenStudy (jdoe0001):

how would you solve say 3(x+4) = x -2 ?

OpenStudy (anonymous):

3*4=7 7-5?

OpenStudy (anonymous):

im bad at math please help

OpenStudy (jdoe0001):

3*4 = 7?

OpenStudy (jdoe0001):

the 3 is multiplying the terms inside, and you'd need to expand

OpenStudy (anonymous):

c?